So I am supposed to do this problem "A 36 gram sample of water is decomposed in an electrolysis reaction giving hydrogen and oxygen. How many grams of hydrogen are produced?" but I am confused on all the converting from grams to moles and moles to grams and am not sure how to get the answer that I want. I know I also have to use the molar ratios somehow but I don't know the exact steps to take.
1
Expert's answer
2019-04-10T08:11:11-0400

During the electrolysis of water the following processes occur

Cathode (-) 2 +2 e=

Anode (+) 4O - 4e= +2O

Total electrolysis equation will be

2 O= 2 +

Find the amount of substance of water

. n (O)= = = 2 mol

Mr (O) = 2Ar (H) +Ar (O) =2×1 +16=18

.n () = n (O) = 2 mol

.m () = n Mr () = 2×2=4 g

Mr () = 2 Ar (H) = 2×1=2

Answer .m () = 4 g